OpenStudy (anonymous):

giving medals and fanning please help What are the solutions to the equation x2 + 4 = 148? A. –4 and 4 B. –12 and 12 C. –2 and 2 D. –11 and 11

OpenStudy (anonymous):

First step is to rearrange the equation, getting everything on one side x^2 - 144 = 0 At this point it's pretty clear what to do Difference of Perfect Squares (x+12)(x-12) = 0 (x+12) = 0, (x-12) = 0 x = -12, 12 B

OpenStudy (anonymous):

thank you so much god bless :)